# Improving LLM-as-a-Judge Inference with the Judgment Distribution

### v1

@@ Abstract
Using language models to scalably approximate human preferences on text quality (LLM-as-a-judge) has become a standard practice applicable to many tasks. A judgment is often extracted from the judge’s textual output alone, typically with greedy decoding. However, LLM judges naturally provide distributions over judgment tokens, inviting a breadth of inference methods for extracting fine-grained preferences. We find that taking the mean of the judgment distribution consistently outperforms taking the mode (i.e. greedy decoding) in all evaluation settings (i.e. pointwise, pairwise, and listwise). We further explore novel methods of deriving preferences from judgment distributions, and find that methods incorporating risk aversion often improve performance. Lastly, we analyze LLM-as-a-judge paired with chain-of-thought (CoT) prompting, showing that CoT can collapse the spread of the judgment distribution, often harming performance. Our findings suggest leveraging distributional output can improve LLM-as-a-judge, as opposed to using the text interface alone.
@@ 1 Introduction
We show that the mean consistently outperforms the mode in the pointwise, pairwise, and listwise settings (i.e. evaluating one, two, and many responses at a time). Specifically, the mean achieves higher accuracy in 92 out of 120 cases on RewardBench Lambert et al. (2024) and MT-Bench Zheng et al. (2023b). We further explore novel methods of deriving preferences from score distributions (Section 4). For example, incorporating risk aversion often improves performance. Categorizing methods as discrete or continuous, where discrete methods (e.g. mode) are simple to interpret like rubric scores, we find that continuous methods outperform discrete methods, due to the latter often predicting ties and failing to capture slight preferences. In particular, the mode assigns ties more frequently than every other method, leading to the lowest accuracy even among discrete methods.
@@ 2.1 LLM-as-a-Judge Settings
Table 1: Mode vs. mean and CoT vs. no-CoT (comma-separated) accuracy results (%). For each base model+setting, we bold the best result and underline results not significantly worse ($\alpha=0.05$). The mean outperforms the mode in 92 out of 120 cases. No-CoT outperforms CoT in 30 out of 40 cases when using the mean for pointwise or pairwise scoring.
@@ 2.2 Related Work
Lee et al. (2024a); Zhai et al. (2024) use pairwise judgment distributions to train a student model, but do not empirically compare with distillation using text judgments. In this work, we benchmark the mode, the mean, and newly proposed methods for leveraging distributional judgments across the pointwise, pairwise, and listwise settings.
@@ 3.2 Experimental Setup
As LLM judges, we use gpt-4o-2024-08-06 (shortened to GPT-4o) OpenAI et al. (2024), Llama-3.1-8B-Instruct (Llama-3.1-8B) Dubey et al. (2024), Mistral-7B-Instruct-v0.3 (Mistral-7B) Jiang et al. (2023), and Prometheus-2-7B Kim et al. (2024). We cover a commonly used closed-sourced LLM (GPT-4o), as well as smaller open-sourced variants.
@@
3.3 Results
Mean outperforms mode
The mean outperforms the mode in 92 out of 120 cases. We observe particularly large gains for pointwise scoring on the Reasoning subset, e.g. absolute +7.7% and +17.1% for GPT-4o and Llama-3.1-8B.
CoT often harms LLM-as-a-judge
For the scoring settings, no-CoT outperforms CoT in 30 out of 40 cases when using the mean. For the pairwise ranking setting, CoT outperforms no-CoT, except with GPT-4o on RewardBench.
We interpret the harmful effect of CoT on pointwise scoring with the smaller models as being due to sharpening, whereby the initial entropy in the judgment is lost as the model commits to one instantiation of a reasoning trace Wang and Zhou (2024). Table 2 confirms this trend by showing that the standard deviation of judgment distributions is lower for CoT than no-CoT. Moreover, removing CoT benefits the mean more than the mode (e.g. 69.6→72.2 for mode vs. 72.7→79.3 for mean, with Llama-3.1-8B on RewardBench), revealing the synergy between eliciting and utilizing distributional judgment.
Which setting works the best?
@@ 3.3 Results
Method RewardBench MT-Bench
Acc $\uparrow$ MSE $\downarrow$ Acc $\uparrow$ MSE $\downarrow$
GPT-4o mode 81.7–${}_{\text{--2.3}}$134+.${}_{\text{+.016}}$ 78.4–${}_{\text{--2.1}}$158+.${}_{\text{+.013}}$
mean 86.7–${}_{\text{--1.3}}$108+.${}_{\text{+.006}}$ 82.9–${}_{\text{--0.3}}$099+.${}_{\text{+.002}}$
[mean] 86.5${}_{\text{+1.3}}$127+.${}_{\text{+.018}}$ 82.7${}_{\text{+2.5}}$182+.${}_{\text{+.036}}$
medi 85.2${}_{\text{+0.6}}$126+.${}_{\text{+.014}}$ 81.5${}_{\text{+1.3}}$170+.${}_{\text{+.028}}$
1p 86.4${}_{\text{+2.1}}$116+.${}_{\text{+.000}}$ 82.7${}_{\text{+1.7}}$165+.${}_{\text{+.027}}$
ram 86.7–${}_{\text{--1.7}}$104+.${}_{\text{+.004}}$ 83.0–${}_{\text{--0.4}}$098+.${}_{\text{+.002}}$
qt 86.6–${}_{\text{--1.3}}$114+.${}_{\text{+.018}}$ 82.7–${}_{\text{--0.5}}$147+.${}_{\text{+.029}}$
ps 86.6–${}_{\text{--1.2}}$105+.${}_{\text{+.009}}$ 82.4–${}_{\text{--0.9}}$107+.${}_{\text{+.004}}$
Llama-3.1-8B mode 72.0–${}_{\text{--0.2}}$221+.${}_{\text{+.029}}$ 75.1${}_{\text{+3.2}}$169+.${}_{\text{+.027}}$
mean 79.3+0.${}_{\text{+0.0}}$156+.${}_{\text{+.001}}$ 81.3–${}_{\text{--0.2}}$103–.${}_{\text{--.001}}$
[mean] 78.5${}_{\text{+3.5}}$198+.${}_{\text{+.012}}$ 80.7${}_{\text{+5.7}}$180+.${}_{\text{+.035}}$
medi 76.5${}_{\text{+2.9}}$207+.${}_{\text{+.016}}$ 80.1${}_{\text{+6.2}}$161+.${}_{\text{+.019}}$
1p 78.5${}_{\text{+2.5}}$195+.${}_{\text{+.012}}$ 81.5${}_{\text{+2.3}}$177+.${}_{\text{+.030}}$
ram 79.7–${}_{\text{--0.2}}$152+.${}_{\text{+.000}}$ 81.1–${}_{\text{--0.3}}$102+.${}_{\text{+.000}}$
qt 78.7–${}_{\text{--0.3}}$177+.${}_{\text{+.013}}$ 81.3${}_{\text{+0.2}}$143+.${}_{\text{+.027}}$
ps 78.6–${}_{\text{--0.3}}$163+.${}_{\text{+.002}}$ 81.8${}_{\text{+0.4}}$111+.${}_{\text{+.001}}$
@@ 4 Study on Pointwise Scoring
We say a method is discrete if it compares two score distributions by their independently assigned scores that take values in $\{1,\ldots,K\}$. Otherwise, we say it is continuous. Discrete scores are often desirable for interpretability (e.g. simple rubrics) but can often result in tied comparisons and fail to capture slight preferences.
@@ 4.1 Methods
Table 3 lists our extended methods for comparing two score distributions. We briefly motivate each newly introduced method below and provide details in Appendix B.1.
We consider the rounded mean as a discrete variant of the mean for a discreteness-controlled comparison with the mode.
1p and ram reflect risk aversion. 1p takes an approach contrary to mode; instead of focusing on where the most mass lies, 1p assigns a low score if there is even a 1% chance of such a low score Siththaranjan et al. (2023). ram is mean but with each distribution shifted down by its risk $\sigma_{-}$.
ps has the intuitive interpretation of the difference in winrates over repeated pairs of samples from the LLM judge Siththaranjan et al. (2023). qt generalizes medi and 1p by averaging the comparisons over all quantiles.
@@ 5.1 Design Decisions
The judge expresses its preference on a $K$-point Likert scale: [>,<][>,<][ > , < ] (Likert-2), [>,=,<][>,=,<][ > , = , < ] (Likert-3), or [,>,=,<,]much-greater-thanmuch-less-than[$\gg$,>,=,<$\ll$][ ≫ , > , = , < , ≪ ] (Likert-5) Liu et al. (2024b).
@@ 5.2 Methods Results
With GPT-4o (not shown; see Tables 14, 15), MSE is always minimized with no-CoT, highlighting the discord between CoT’s sharpening effect and calibration. In Appendix F.3, we further analyze position bias and find that CoT increases the occurrence of severe position bias.
@@ 5.3 Granularity Results
Table 8 compares the Likert scales used in the pairwise ranking prompt. We find that Likert-2 performs the best overall, in line with the AlpacaEval methodology Dubois et al. (2024) but deviating from WB-Reward and Arena-Hard-Auto Lin et al. (2024); Li et al. (2024b), which use Likert-5.222We remark that the evaluation setups in these works and ours differ in that theirs use coarse-grained, model-level agreement using Chatbot Arena Elo scores Zheng et al. (2023b), while ours uses fine-grained, instance-level agreement.
@@ 6 Listwise Judgment
Listwise judgment is not as prevalent as pointwise or pairwise judgment, but is more efficient Zhu et al. (2024). Furthermore, listwise prompting grants the judge the maximal context for comparison Buyl et al. (2023), generalizing the advantage of pairwise prompting over pointwise prompting, given a sufficiently capable judge.
@@ 6.2 Experimental Setup
Due to the context length required for listwise ranking and the difficulty of the task, we limit our evaluation to GPT-4o.
@@ 7 Conclusion and Recommendations
Acknowledgments
We thank the members of the UT NLP community, especially Manya Wadhwa for feedback on the draft.
@@ Appendix B Methods
Let $A_{1}$ and $A_{2}$ be two texts to compare. We describe the methods of predicting a value in $[-1,1]$ that signifies the advantage of $A_{1}$ over $A_{2}$. For accuracy, we take the sign of the prediction. For MSE, we rescale predictions from $[-1,1]$ to $[0,1]$.
The methods are invariant to scaling and translating the judgment space, and all methods that do not take expectations $\mathbb{E}$ (which assumes linearity) are invariant to applying a positive monotone transformation to the judgment space.
@@ B.1 Pointwise Methods
We elaborate on the pointwise methods introduced in Section 4.1. The LLM judge independently judges $A_{1}$ and $A_{2}$, producing score distributions over $\{1,\ldots,K\}$ for an integer $K$ that define independent random variables $X_{1}$ and $X_{2}$, which are used to compare $A_{1}$ and $A_{2}$. The methods are all equivalent if the distributions are deterministic, thus our experiments evaluate their ability to leverage the LLM judge’s distributional output.
The denominator in mean normalizes it into $[-1,1]$, similar to ${\rm sgn}(x)=\frac{x}{|x|}$, taking $\frac{0}{0}$ to be 0. The $\sigma$ term lowers the magnitude of the prediction in the presence of uncertainty in a continuous manner. Specifically, let $k,k^{\prime}\in\{1,\ldots,K\}$ with $k\neq k^{\prime}$. For $\epsilon\in[0,1]$, let $X_{1}$ have a two-point distribution $(1-\epsilon)\delta_{k}+\epsilon\delta_{k^{\prime}}$ and let $X_{2}$ have a deterministic distribution $\delta_{k}$. Then $\textsc{mean}(X_{1},X_{2})$ as a function of $\epsilon$ is continuous at $\epsilon=0$.
@@ B.2 Pairwise Methods
[…]
@@ B.2.1 Pairwise Scoring
[…]
@@
agg-mode :-$\coloneq{\rm sgn}({\rm mode}(X_{\Delta})):$
agg-medi :-$\coloneq{\rm sgn}({\rm median}(X_{\Delta})):$
agg-mean :-$\coloneq\textsc{mean}(X_{\Delta}),:$
@@
mode-agg :-$\coloneq\frac{\sum_{\mathbf{o}\in\mathbf{O}}{\rm mode}(X^{\mathbf{o}}_{\Delta})}{\sum_{\mathbf{o}\in\mathbf{O}}|{\rm mode}(X^{\mathbf{o}}_{\Delta})|}:$
medi-agg :-$\coloneq\frac{\sum_{\mathbf{o}\in\mathbf{O}}{\rm median}(X^{\mathbf{o}}_{\Delta})}{\sum_{\mathbf{o}\in\mathbf{O}}|{\rm median}(X^{\mathbf{o}}_{\Delta})|}:$
mean-agg :-$\coloneq\frac{1}{|\mathbf{O}|}\sum_{\mathbf{o}\in\mathbf{O}}\textsc{mean}(X^{\mathbf{o}}_{\Delta}),:$
taking 00:-0:-${0}\coloneq 0$ :- 0.

### v2

@@ Abstract
Using language models to scalably approximate human preferences on text quality (LLM-as-a-judge) has become a standard practice applicable to many tasks. A judgment is often extracted from the judge’s textual output alone, typically with greedy decoding. However, LLM judges naturally provide distributions over judgment tokens, inviting a breadth of inference methods for extracting fine-grained preferences. We find that taking the mean of the judgment distribution consistently outperforms taking the mode (i.e. greedy decoding) in all evaluation settings (i.e. pointwise, pairwise, and listwise). We further explore novel methods of deriving preferences from judgment distributions, and find that methods incorporating risk aversion often improve performance. Lastly, we analyze LLM-as-a-judge paired with chain-of-thought (CoT) prompting, showing that CoT can collapse the spread of the judgment distribution, often harming performance. Our findings show that leveraging distributional output improves LLM-as-a-judge, as opposed to using the text interface alone.
@@ 1 Introduction
We show that the mean consistently outperforms the mode in the pointwise, pairwise, and listwise settings (i.e. evaluating one, two, and many responses at a time). Specifically, the mean achieves higher accuracy in 42 out of 48 cases on RewardBench Lambert et al. (2024) and MT-Bench Zheng et al. (2023b). We further explore novel methods of deriving preferences from score distributions (Section 4). For example, incorporating risk aversion often improves performance. Categorizing methods as discrete or continuous, where discrete methods (e.g. mode) are simple to interpret like rubric scores, we find that continuous methods outperform discrete methods, due to the latter often predicting ties and failing to capture slight preferences. In particular, the mode assigns ties more frequently than every other method, leading to the lowest accuracy even among discrete methods.
@@ 2.2 Related Work
Lee et al. (2024a); Zhai et al. (2024) use pairwise judgment distributions to train a student model, but do not empirically compare with distillation using one-hot judgments. In this work, we benchmark the mode, the mean, and newly proposed methods for leveraging distributional judgments across the pointwise, pairwise, and listwise settings.
@@ 3.2 Experimental Setup
As LLM judges, we use gpt-4o-2024-08-06 (shortened to GPT-4o) OpenAI et al. (2024), Llama-3.1-8B-Instruct (Llama-3.1-8B) Dubey et al. (2024), Mistral-7B-Instruct-v0.3 (Mistral-7B) Jiang et al. (2023), and Prometheus-2-7B Kim et al. (2024). We cover a commonly used closed-source LLM [^2] Some proprietary LLMs such as Anthropic Claude do not provide logit access, preventing us from including them in our experiments. In Appendix E.1, we provide partial results for DeepSeek-V3, whose trends match those of GPT-4o. (GPT-4o), as well as smaller open-source variants.
@@
Model Setting Method Reward MT-Bench Bench GPT-4o point score mode 85.1, 84.0 81.9, 80.5 mean 87.4, 88.0 83.6, 83.2 pair score mode 86.7, 87.4 86.2, 86.5 mean 87.1, 87.6 86.3, 86.8 pair rank mode 88.4, 89.7 86.3, 85.6 mean 88.6, 90.5 87.3, 85.9 Llama -3.1-8B point score mode 69.6, 72.2 74.9, 71.9 mean 72.7, 79.3 78.7, 81.5 pair score mode 71.7, 75.2 82.6, 82.4 mean 72.1, 76.8 82.3, 81.2 pair rank mode 68.9, 58.9 76.2, 63.0 mean 74.2, 68.6 80.0, 76.5 Mistral-7B point score mode 60.4, 62.7 59.5, 66.2 mean 63.8, 72.1 62.6, 74.0 pair score mode 67.3, 68.9 79.3, 79.8 mean 68.1, 71.0 80.0, 80.4 pair rank mode 56.3, 53.8 51.5, 51.5 mean 63.9, 59.1 73.5, 65.5 Prometheus -2-7B point score mode 64.3, 66.0 72.5, 73.5 mean 64.6, 75.2 72.1, 81.6 pair score mode 71.0, 68.7 78.4, 80.8 mean 70.5, 70.8 78.3, 80.9 pair rank mode 59.6, 48.2 51.5, 43.0 mean 69.7, 48.8 75.4, 33.4
Table 1: Mode vs. mean and CoT vs. no-CoT (comma-separated) accuracy results (%). For each base model+setting, we bold the best result and underline results not significantly worse ($\alpha=0.05$). The mean outperforms the mode in 42 out of 48 cases. No-CoT outperforms CoT in 14 out of 16 cases when using the mean for pointwise or pairwise scoring.
@@
3.3 Results
Mean outperforms mode
The mean outperforms the mode in 42 out of 48 cases. In Table 10, we provide a subset breakdown of RewardBench and observe particularly large gains for pointwise scoring on the Reasoning subset.
@@ 3.3 Results
CoT often harms LLM-as-a-judge
For the scoring settings, no-CoT outperforms CoT in 14 out of 16 cases when using the mean. For the pairwise ranking setting, CoT outperforms no-CoT, except with GPT-4o on RewardBench.
We interpret the harmful effect of CoT on pointwise scoring with the smaller models as being due to sharpening, whereby the initial entropy in the judgment is lost as the model commits to one instantiation of a reasoning trace. Table 2 confirms this trend by showing that the standard deviation of judgment distributions is lower for CoT than no-CoT. Moreover, removing CoT benefits the mean more than the mode (e.g. 69.6→72.2 for mode vs. 72.7→79.3 for mean, with Llama-3.1-8B on RewardBench), revealing the synergy between eliciting and utilizing distributional judgment.
Which setting works the best?
@@ 4 Study on Pointwise Scoring
We say a method is discrete if it compares two score distributions by their independently assigned scores that take values in $\{1,\ldots,K\}$. Otherwise, we say it is continuous. Discrete scores are often desirable for interpretability (e.g. simple rubrics) but, by the pigeonhole principle, can often result in tied comparisons and fail to capture slight preferences.
@@ 4.1 Methods
Table 3 lists our extended methods for comparing two score distributions. We motivate the newly introduced methods below and provide details in Appendix B.1.
Users often prefer discrete methods (e.g. mode) because they are simple to interpret, even if they have lower accuracy than continuous methods (e.g. mean). This motivates the question of where the mode (the status quo method) ranks among discrete methods. To answer this question, we compare the mode to other discrete methods: rounded mean, median, and first percentile (discussed in the next paragraph).
Humans exhibit risk aversion when making decisions. They often disprefer negative outcomes more strongly than they prefer positive outcomes Holt and Laury (2002). However, this disposition is not captured by the measures of central tendency discussed so far. Thus, we investigate whether incorporating the human disposition of risk aversion into LLM-as-a-judge inference methods improves alignment with human preferences. The methods 1p (discrete) and ram (continuous) reflect risk aversion. 1p takes an approach contrary to mode; instead of focusing on where the most mass lies, 1p assigns a low score if there is even a 1% chance of such a low score Siththaranjan et al. (2023). ram is mean but with each distribution shifted down by its risk $\sigma_{-}$.
We have used mode to represent the status quo LLM-as-a-judge inference method, which uses greedy decoding to obtain a judgment token. However, some prior works use a positive temperature, e.g. to obtain varied CoT chains Zhang et al. (2024a), in which case a sampled judgment token is decoded rather than the mode. To account for the random nature of sampling, we design the method ps as the difference in winrates over repeated pairs of samples from the LLM judge Siththaranjan et al. (2023). qt generalizes medi and 1p by averaging the comparisons over all quantiles, and can be viewed as ps but with $X_{1}$ and $X_{2}$ positively monotonically correlated.
@@ 4.2 Results
Method RewardBench MT-Bench
Acc $\uparrow$ MSE $\downarrow$ Acc $\uparrow$ MSE $\downarrow$
GPT-4o
[…]
Llama-3.1-8B
[…]
@@ 5.1 Design Decisions
We prompt the judge to express its preference on a $K$-point Likert scale: $[>,<]$ (Likert-2), $[>,=,<]$ (Likert-3), or [,>,=,<,][$\gg$$,\ll$] (Likert-5) Liu et al. (2024b). In the prompt (Appendix C.3), the symbols are assigned descriptions such as “significantly better” and “slightly better”.
@@ 5.2 Methods Results
With GPT-4o (not shown; see Tables 15, 16), MSE is always minimized with no-CoT, highlighting the discord between CoT’s sharpening effect and calibration. In Appendix F.3, we further analyze position bias and find that CoT increases the occurrence of severe position bias.
@@ 5.3 Granularity Results
Table 8 compares the Likert scales used in the pairwise ranking prompt. We find that Likert-2 performs the best overall, in line with the AlpacaEval methodology Dubois et al. (2024) but deviating from WB-Reward and Arena-Hard-Auto Lin et al. (2024); Li et al. (2024b), which use Likert-5.
@@ 6 Listwise Judgment
Listwise judgment is not as prevalent as pointwise or pairwise judgment, but it offers efficiency Zhu et al. (2024) while granting the judge the maximal context for comparison Buyl et al. (2023).
@@ 6.2 Experimental Setup
Due to the context length required for listwise ranking and the difficulty of the task, we limit our evaluation to GPT-4o. In preliminary experiments, we found poor performance with the smaller models, but in Appendix E.1 we show that DeepSeek-V3 exhibits similar trends to GPT-4o.
@@ Appendix B Methods
Let $A_{1}$ and $A_{2}$ be two texts to compare. We describe the methods of predicting a value in [1,1][-1,1] that signifies the advantage of $A_{1}$ over $A_{2}$. For accuracy, we take the sign of the prediction. For MSE, we rescale predictions from [1,1][-1,1] to $[0,1]$.
@@ B.1 Pointwise Methods
We elaborate on the pointwise methods introduced in Section 4.1. The LLM judge independently judges $A_{1}$ and $A_{2}$, producing score distributions over $\{1,\ldots,K\}$ for an integer $K$ that define independent random variables $X_{1}$ and $X_{2}$, which are used to compare $A_{1}$ and $A_{2}$.
The methods are invariant to scaling and translating the judgment space, and all methods that do not take expectations $\mathbb{E}$ (which assumes linearity) are invariant to applying a positive monotone transformation to the judgment space. The methods are all equivalent if the distributions are deterministic, thus our experiments evaluate their ability to leverage the LLM judge’s distributional output.
The denominator in mean normalizes it into [1,1][-1,1], similar to ${\rm sgn}(x)=\frac{x}{|x|}$, taking $\frac{0}{0}$ to be 0. The $\sigma$ term lowers the magnitude of the prediction in the presence of uncertainty in a continuous manner. Specifically, let $k,k^{\prime}\in\{1,\ldots,K\}$ with $k\neq k^{\prime}$. For $\epsilon\in[0,1]$, let $X_{1}$ have a two-point distribution $(1-\epsilon)\delta_{k}+\epsilon\delta_{k^{\prime}}$ and let $X_{2}$ have a deterministic distribution $\delta_{k}$. Then $\textsc{mean}(X_{1},X_{2})$ as a function of $\epsilon$ is continuous at $\epsilon=0$.
@@ B.2 Pairwise Methods
To account for position bias, we prompt the LLM judge once for each order of presentation. For an order $\mathbf{o}\in\mathbf{O}\coloneq\{(1,2),(2,1)\}$, we use $\mathbf{o}$ to denote dependence on the order $(A_{o_{1}},A_{o_{2}})$ in which the texts appear in the prompt.
@@ B.2.1 Pairwise Scoring
$X^{\mathbf{o}}_{\Delta}\overset{d}{=}(X^{\mathbf{o}}_{1}-X^{\mathbf{o}}_{2})\mid(X^{\mathbf{o}}_{o_{1}}=x_{o_{1}})$
@@
agg-mode $\coloneq{\rm sgn}({\rm mode}(X_{\Delta}))$
agg-medi $\coloneq{\rm sgn}({\rm median}(X_{\Delta}))$
agg-mean $\coloneq\textsc{mean}(X_{\Delta}),$
@@
mode-agg $\coloneq\frac{\sum_{\mathbf{o}\in\mathbf{O}}{\rm mode}(X^{\mathbf{o}}_{\Delta})}{\sum_{\mathbf{o}\in\mathbf{O}}|{\rm mode}(X^{\mathbf{o}}_{\Delta})|}$
medi-agg $\coloneq\frac{\sum_{\mathbf{o}\in\mathbf{O}}{\rm median}(X^{\mathbf{o}}_{\Delta})}{\sum_{\mathbf{o}\in\mathbf{O}}|{\rm median}(X^{\mathbf{o}}_{\Delta})|}$
mean-agg $\coloneq\frac{1}{|\mathbf{O}|}\sum_{\mathbf{o}\in\mathbf{O}}\textsc{mean}(X^{\mathbf{o}}_{\Delta}),$
taking $\frac{0}{0}\coloneq 0$.